OpenStudy (anonymous):

Alex stocks up for winter. He buys 32 cans of vegetables. He pays 80 cents per can for tomatoes and 40 cents per can for corn, for a total cost of $18.00. How many cans of corn does he buy?

OpenStudy (anonymous):

I've got 80x+40y=1800 And 80x+40y=32

OpenStudy (texaschic101):

80x + 40y = 1800 x + y = 32

OpenStudy (anonymous):

Yes, it's plug and go after that, how would I go from there, though?

OpenStudy (anonymous):

x=13, y=19.

OpenStudy (anonymous):

I'm wrong still.

OpenStudy (texaschic101):

x + y = 32 --> x = 32 - y now sub that into the other equation 80(32 - y) + 40y = 1800 does that help....now you solve for y. Once y is found, sub that back into the equation x + y = 32, and solve for x if you wish to check your answer, sub it back into the equations.. x + y = 32 13 + 19 = 32 32 = 32 (this works) 80x + 40y = 1800 80(13) + 40(19) = 1800 1040 + 760 = 1800 1800 = 1800 (this works) you are correct..x = 13 and y = 19

OpenStudy (anonymous):

I saw that when I plugged it in. Thank you for the steps. I have a similar question so this should help.

OpenStudy (texaschic101):

glad I could help :)